OpenStudy (anonymous):

A population of 125 bacteria triples in size every 5 hours. Use the function f(x) = 125 • 3x to find the number of bacteria after 15 hours. Let x = the number of 5 hour time periods. a. 3375 b. 3054 c. 2261 d. 1793

OpenStudy (anonymous):

After 5 hours you have 125*3=375 After 10 hours it is 375*3=1125 After 15 hours its 1125*3=3375

OpenStudy (anonymous):

So it's A right?

OpenStudy (anonymous):

yes